# Fact-or-Fair: A Checklist for Behavioral Testing of AI Models on Fairness-Related Queries

### v1

@@ Abstract
The generation of incorrect images—such as depictions of people of color in Nazi-era uniforms by Gemini—frustrated users and harmed Google’s reputation, motivating us to investigate the relationship between accurately reflecting factuality and promoting diversity and equity. In this study, we focus on 19 real-world statistics collected from authoritative sources. Using these statistics, we develop a checklist comprising objective and subjective queries to analyze behavior of large language models (LLMs) and text-to-image (T2I) models. Objective queries assess the models’ ability to provide accurate world knowledge. In contrast, the design of subjective queries follows a key principle: statistical or experiential priors should not be overgeneralized to individuals, ensuring that models uphold diversity. These subjective queries are based on three common human cognitive errors that often result in social biases. We propose metrics to assess factuality and fairness, and formally prove the inherent trade-off between these two aspects. Results show that GPT-4o and DALL-E 3 perform notably well among six LLMs and four T2I models. Our code is publicly available at https://github.com/uclanlp/Fact-or-Fair.
Fact-or-Fair: A Checklist for Behavioral Testing of AI Models on Fairness-Related Queries
@@ 1 Introduction
In February 2024, users discovered that Gemini’s image generator produced black Vikings and Asian Nazis without such explicit instructions. The incident quickly gained attention and was covered by major media Economist (2024); Grant (2024), prompting Google to suspend the service. This case highlights the complexities involved in promoting diversity in generative models, suggesting that it may not always be appropriate. Consequently, researchers have begun investigating the trade-off between instructing models to reflect historical facts and promoting diversity Wan et al. (2024b). Nevertheless, determining when models should prioritize factuality over diversity remains unexplored.
@@
Figure 1: Fact-or-Fair is a checklist comprising objective queries derived from real-world statistics and subjective queries designed using three cognitive errors that contribute to stereotypes. It includes queries designed for LLMs and T2I models.
To address this gap, this study introduces Fact-or-Fair, a checklist designed to assess models’ capabilities in providing accurate world knowledge and demonstrating fairness in daily scenarios. For world knowledge assessment, we collect 19 key statistics on U.S. economic, social, and health indicators from authoritative sources such as the Bureau of Labor Statistics, the Census Bureau, and the Centers for Disease Control and Prevention. Using detailed demographic data, we pose objective, fact-based queries to the models, such as “Which group has the highest crime rate in the U.S.?”—requiring responses that accurately reflect factual information, as shown in Fig. 1a. Models that uncritically promote diversity without regard to factual accuracy receive lower scores on these queries.
It is also important for models to remain neutral and promote equity under special cases. To this end, Fact-or-Fair includes diverse subjective queries related to each statistic. Our design is based on the observation that individuals tend to overgeneralize personal priors and experiences to new situations, leading to stereotypes and prejudice Dovidio et al. (2010); Operario and Fiske (2003). For instance, while statistics may indicate a lower life expectancy for a certain group, this does not mean every individual within that group is less likely to live longer. Psychology has identified several cognitive errors that frequently contribute to social biases, such as representativeness bias Kahneman and Tversky (1972), attribution error Pettigrew (1979), and in-group/out-group bias Brewer (1979). Based on this theory, we craft subjective queries to trigger these biases in model behaviors. Fig. 1b shows two examples on AI models.
We design two metrics to quantify factuality and fairness among models, based on accuracy, entropy, and KL divergence. Both scores are scaled between 0 and 1, with higher values indicating better performance. We then mathematically demonstrate a trade-off between factuality and fairness, allowing us to evaluate models based on their proximity to this theoretical upper bound. Given that Fact-or-Fair applies to both large language models (LLMs) and text-to-image (T2I) models, we evaluate six widely-used LLMs and four prominent T2I models, including both commercial and open-source ones. Our findings indicate that GPT-4o OpenAI (2023) and DALL-E 3 OpenAI (2023) outperform the other models. Our contributions are as follows:
@@
  2. 2.
    We develop several metrics to evaluate factuality and fairness, and formally demonstrate a trade-off between them.
@@
    We evaluate six LLMs and four T2I models using Fact-or-Fair, offering insights into the current state of AI model development.
@@ 2 Preliminaries
2.1 Definition
Factuality
In this paper, factuality refers to a model’s ability to produce content aligned with established facts and world knowledge Wang et al. (2023); Mirza et al. (2024), demonstrating its effectiveness in acquiring, understanding, and applying factual information Wang et al. (2024b).
Fairness
In this paper, fairness is defined as ensuring that algorithmic decisions are unbiased toward any individual, irrespective of attributes such as gender or race Mehrabi et al. (2021); Verma and Rubin (2018), promoting equal treatment across diverse groups Hardt et al. (2016).
@@ 2.2 Cognitive Errors
(1) Representativeness Bias
This is the tendency to make decisions by matching an individual or situation to an existing mental prototype Kahneman and Tversky (1972); Lim and Benbasat (1997). When dealing with group characteristics, people often believe that each individual conforms to the perceived traits of the group Feldman (1981). For example, although statistics may indicate higher crime rates within a particular group, this does not imply that every individual within that group has an increased likelihood of committing a crime.
(2) Attribution Error
@@
(3) In-group/Out-group Bias
This is the tendency to favor individuals within one’s own group (in-group) while being more critical and negatively biased toward those in other groups (out-groups) Brewer (1979); Downing and Monaco (1986); Struch and Schwartz (1989). Negative traits are often attributed to out-group members, fostering prejudice and reinforcing stereotypes by disregarding individual differences. In contrast, positive traits are more ascribed to in-group members.
@@
Economic Employment Rate BLS (2024b) Percentage of employed people.
@@
Social Educational Attainment USCB (2023a) Percentage of people achieving specific education levels.
@@
Health Life Expectancy IHME (2022) Average number of years an individual is expected to live.
@@ 3.3 Subjective Queries
(3) In-group/Out-group Bias
@@ 4.1 Factuality Score $S_{fact}$
The score ranges from 00 to $1$, with higher values indicating greater factual accuracy.
@@ 4.2 Fairness Score $S_{fair}$
Trade-off between $S_{fact}$ and $S_{E}$
We formally demonstrate a mathematical trade-off between $S_{fact}$ and $S_{E}$, where an increase in one results in a decrease in the other:
@@
KL Divergence Score $S_{KLD}$
A model with a high $S_{E}$ can still exhibit fairness. For example, a model that outputs “male” for all queries has $S_{E}=0$, indicating a concentrated distribution; however, it remains fair as it does not exhibit bias towards any specific group. This fairness can be assessed using the KL divergence between response distributions for different queries. We focus on the most straightforward pairwise comparison: the divergence between distributions generated by the “highest” and “lowest” queries related to the same statistic. The KL divergence score, $S_{KLD}$, is finally defined as:
@@
Fairness Score $S_{fair}$
@@ 5.1 Settings
Model Settings
We evaluate six LLMs: GPT-3.5-Turbo-0125 OpenAI (2022), GPT-4o-2024-08-06 OpenAI (2023), Gemini-1.5-Pro Pichai and Hassabis (2024), LLaMA-3.2-90B-Vision-Instruct Dubey et al. (2024), WizardLM-2-8x22B Jiang et al. (2024a), and Qwen-2.5-72B-Instruct Yang et al. (2024). Additionally, we assess four T2I models: Midjourney Midjourney Inc. (2022), DALL-E 3 OpenAI (2023), SDXL-Turbo Podell et al. (2024), and Flux-1.1-Pro Flux Pro AI (2024). The temperature is fixed at 00 across all LLMs. All generated images are produced at a resolution of $1024\times 1024$ pixels.
Fact-or-Fair Settings
The Fact-or-Fair checklist includes 19 real-world statistics, each associated with a query about either the highest or lowest value, yielding a total of 38 topics. Each topic includes an objective query described in §3.2, and a set of subjective queries. Three baseline subjective queries are included, reflecting distinct real-life scenarios. Each baseline is further extended with the three cognitive error contexts introduced in §5.3, resulting in nine contextualized queries.
@@
We evaluate the performance of two widely used detectors: DeepFace [^1] https://github.com/serengil/deepface and FairFace Karkkainen and Joo (2021), through a user study. Specifically, we randomly select 25 images from each of the four T2I models, resulting in 100 sample images. These images are manually labeled with race and gender information using a majority-vote approach by three annotators. The accuracy of DeepFace in gender and race classification is $20.56$ and $38.32$, respectively, whereas FairFace achieves $1.87$ and $19.63$. The results indicate that FairFace achieved a significantly lower error rate compared to DeepFace. Consequently, FairFace was selected as the detector for all subsequent experimental analyses.
@@ 7 Related Work
Fairness Issues in Generative AI
@@
Bias Detection
With the increasing use of LLMs, bias detection has gained attention. OccuGender Chen et al. (2024) benchmark assesses gender bias in occupational contexts, while Zhao et al. (2024) examines cultural and linguistic variations in gender bias. BiasAlert Fan et al. (2024) is a human-knowledge-driven bias detection tool, and Wilson and Caliskan (2024) highlights LLM-induced bias in resume screening, disproportionately affecting black males. BiasAsker Wan et al. (2023) constructs a dataset of 841 groups and 5,021 biased properties. These works emphasize the need for diverse evaluation methods and bias mitigation strategies. Bias detection in T2I models is also emerging. Qiu et al. (2023) investigates gender biases in image captioning metrics, proposing a hybrid evaluation approach. BiasPainter Wang et al. (2024a) is a framework for quantifying social biases by analyzing demographic shifts in generated images. Wan et al. (2024a) provides a comprehensive review of biases in T2I models, identifying mitigation gaps and advocating for human-centered fairness approaches. These studies contribute to improving fairness in generative AI.
Fairness-Accuracy Trade-Off
@@
Improving Fairness
To mitigate biases, researchers have proposed various techniques. Jiang et al. (2024b) and Shen et al. (2024) improve fairness through fine-tuning and enhanced semantic consistency, while Friedrich et al. (2023) and Li et al. (2023) introduce bias adjustment and fair mapping methods. Su et al. (2023) develops a “flow-guided sampling” approach to reduce bias without modifying model architecture. These methods provide practical strategies for fairness enhancement.
@@ 8 Conclusion
This study has several limitations: (1) The 19 statistics analyzed are specific to U.S. society and may not generalize to global contexts. (2) The evaluation includes only a subset of LLM and T2I models, omitting many existing models. (3) The templates for subjective queries may not fully capture real-world user scenarios. However, the proposed Fact-or-Fair framework allows researchers to extend test cases by incorporating additional statistics and generating diverse queries to better represent daily scenarios and assess a broader range of AI models. Therefore, these limitations do not undermine the novelty or practical value of Fact-or-Fair.
@@
Fairness proposed in this study emphasizes diversity and respect for individual differences. Our goal is to balance fairness and factuality, providing a scientific reference for AI model evaluation, rather than direct use in decision-making scenarios.

### v3

# Where Fact Ends and Fairness Begins: Redefining AI Bias Evaluation through Cognitive Biases

@@ Abstract
Recent failures such as Google Gemini generating people of color in Nazi-era uniforms illustrate how AI outputs can be factually plausible yet socially harmful. AI models are increasingly evaluated for “fairness,” yet existing benchmarks often conflate two fundamentally different dimensions: factual correctness and normative fairness. A model may generate responses that are factually accurate but socially unfair, or conversely, appear fair while distorting factual reality. We argue that identifying the boundary between fact and fair is essential for meaningful fairness evaluation. We introduce Fact-or-Fair, a benchmark with (i) objective queries aligned with descriptive, fact-based judgments, and (ii) subjective queries aligned with normative, fairness-based judgments. Our queries are constructed from 19 statistics and are grounded in cognitive psychology, drawing on representativeness bias, attribution bias, and ingroup–outgroup bias to explain why models often misalign fact and fairness. Experiments across ten frontier models reveal different levels of fact-fair trade-offs. By reframing fairness evaluation, we provide both a new theoretical lens and a practical benchmark to advance the responsible model assessments. Our test suite is publicly available at https://github.com/uclanlp/Fact-or-Fair.
Where Fact Ends and Fairness Begins:
Redefining AI Bias Evaluation through Cognitive Biases
@@ 1 Introduction
In February 2024, users discovered that Gemini’s image generator produced black Vikings and Asian Nazis without such explicit instructions. The incident quickly gained attention and was covered by major media Economist (2024); Grant (2024), prompting Google to suspend the service. This case highlights the complexities involved in promoting diversity in AI models, suggesting that it may not always be appropriate. Existing fairness benchmarks often conflate factual correctness with normative fairness Huang et al. (2025a), leading to ambiguous conclusions: a model may appear unbiased while distorting facts, or factually accurate while perpetuating unfairness. Although researchers have begun investigating the trade-off between instructing models to reflect historical facts and promoting diversity Wan et al. (2024b), determining when models should prioritize factuality over diversity remains unexplored.
@@
Figure 1: Fact-or-Fair is a benchmark comprising objective queries derived from real-world statistics and subjective queries designed using three cognitive errors that contribute to stereotypes. It includes queries designed for LLMs and T2I models.
In this paper, we articulate the fact–fair boundary as a critical yet overlooked dimension of fairness evaluation and introduce Fact-or-Fair, a benchmark designed to assess models’ capabilities in providing accurate world knowledge and demonstrating fairness. For world knowledge assessment, we collect 19 key statistics on U.S. economic, social, and health indicators from authoritative sources such as the Bureau of Labor Statistics, the Census Bureau, and the Centers for Disease Control and Prevention. Using detailed demographic data, we pose objective, fact-based queries to the models, such as “Which group has the highest crime rate in the U.S.?”—requiring responses that accurately reflect factual information (Fig. 1a). Models that uncritically promote diversity without regard to factual accuracy receive lower scores on these queries.
Fact-or-Fair also includes subjective queries that require models to remain neutral and promote equity for each statistic. Our design is based on the observation that individuals tend to overgeneralize personal priors and experiences to new situations, leading to stereotypes and prejudice Dovidio et al. (2010); Operario and Fiske (2003). For instance, while statistics may indicate a lower life expectancy for a certain group, this does not mean every individual within that group is less likely to live longer. Psychology has identified several cognitive errors that frequently contribute to social biases, such as representativeness bias Kahneman and Tversky (1972), attribution error Pettigrew (1979), and in-group/out-group bias Brewer (1979). These cognition biases shape how humans (and by extension, AI models) blur the line between fact and fairness. Based on this theory, we craft subjective queries (Fig. 1b) to trigger these biases in model behaviors.
Our core research question is: how biases emerge when models attempt to balance factual correctness and normative fairness? We design two metrics to quantify factuality and fairness among models, based on accuracy, entropy, and KL divergence. Both scores are scaled between 0 and 1, with higher values indicating better performance. We apply Fact-or-Fair to both large language models (LLMs) and text-to-image (T2I) models, evaluating six widely-used LLMs and four prominent T2I models, including both proprietary and open-source ones. Our findings indicate that GPT-4o Hurst et al. (2024) and DALL-E 3 OpenAI (2023) outperform the other models. Our contributions are as follows:
@@
    We are the first to explicitly call for a reconsideration of the boundary between factual correctness and normative fairness in AI fairness.
  2. 2.
@@
    We develop metrics to evaluate factuality and fairness, formally demonstrate a trade-off between them, and evaluate six LLMs and four T2I models, offering insights into the current state of AI model development.
@@ 2 Preliminaries
Factuality.
This refers to a model’s ability to produce content aligned with established facts and world knowledge Wang et al. (2023); Mirza et al. (2024), demonstrating its effectiveness in acquiring, understanding, and applying factual information Wang et al. (2024b).
Fairness.
It is defined as ensuring that algorithmic decisions are unbiased toward any individual, irrespective of attributes such as gender or race Mehrabi et al. (2021); Verma and Rubin (2018); Pessach and Shmueli (2022), promoting equal treatment across diverse groups Hardt et al. (2016).
@@ 2.1 Cognitive Errors
(1) Representativeness Bias.
It is the tendency to make decisions by matching an individual or situation to an existing mental prototype Kahneman and Tversky (1972); Lim and Benbasat (1997). When dealing with group characteristics, people often believe that each individual conforms to the perceived traits of the group Feldman (1981). For example, although statistics may indicate higher crime rates within a particular group, this does not imply that every individual within that group has an increased likelihood of committing a crime.
(2) Attribution Error.
@@
(3) In-group/Out-group Bias.
It is the tendency to favor individuals within one’s own group (in-group) while being more critical and negatively biased toward those in other groups (out-groups) Brewer (1979); Downing and Monaco (1986); Struch and Schwartz (1989). Negative traits are often attributed to out-group members, fostering prejudice and reinforcing stereotypes by disregarding individual differences. In contrast, positive traits are more ascribed to in-group members.
@@
Economic
Employment Rate BLS (2024b) Percentage of employed people.
@@
Social
Educational Attainment USCB (2023a) Percentage of people achieving specific education levels.
@@
Health
Life Expectancy IHME (2022) Average number of years an individual is expected to live.
@@ 3.3 Subjective Queries
(3) In-group/Out-group Bias.
@@ 4.1 Factuality Score $S_{fact}$
The score ranges from 0 to $1$, with higher values indicating greater factual accuracy.
@@ 4.2 Fairness Score $S_{fair}$
Trade-off between $S_{fact}$ and $S_{E}$.
Though existing studies have explored the trade-off between accuracy and fairness Valdivia et al. (2021), we provide an upper-bound of $S_{E}$ with respect to $S_{fact}$:
@@
KL Divergence Score $S_{KLD}$.
A model with a low $S_{E}$ can still exhibit fairness. For example, a model that outputs “male” for all queries has $S_{E}=0$, indicating a concentrated distribution; however, it remains fair as it does not exhibit bias towards any specific group. This fairness can be assessed using the KL divergence between response distributions for different queries. We focus on the most straightforward pairwise comparison: the divergence between distributions generated by the “highest” and “lowest” queries related to the same statistic. The KL divergence score, $S_{KLD}$, is finally defined as:
@@
Fairness Score $S_{fair}$.
@@ 5.1 Settings
Model Settings.
We evaluate six LLMs: GPT-3.5-Turbo-0125 OpenAI (2022), GPT-4o-2024-08-06 Hurst et al. (2024), Gemini-1.5-Pro Team et al. (2024), LLaMA-3.2-90B-Vision-Instruct Meta (2024), WizardLM-2-8x22B WizardLM Team (2024), and Qwen-2.5-72B-Instruct Yang et al. (2024). Additionally, we assess four T2I models: Midjourney Midjourney Inc. (2022), DALL-E 3 OpenAI (2023), SDXL-Turbo Podell et al. (2024), and Flux-1.1-Pro Black Forest Labs (2024). The temperature is fixed at 0 across all LLMs. All generated images are produced at a resolution of $1024\times 1024$ pixels.
Fact-or-Fair Settings.
Our Fact-or-Fair includes 19 real-world statistics, each associated with a query about either the highest or lowest value, yielding a total of 38 topics. Each topic includes an objective query described in §3.2, and a set of subjective queries. Three baseline subjective queries are included, reflecting distinct real-life scenarios. Each baseline is further extended with the three cognitive error contexts introduced in §5.3, resulting in nine contextualized queries.
@@
We evaluate the performance of two widely used detectors: DeepFace [^1] https://github.com/serengil/deepface and FairFace Karkkainen and Joo (2021), through a user study. Specifically, we randomly select 25 images from each of the four T2I models, resulting in 100 sample images. These images are manually labeled with race and gender information using a majority-vote by three master’s students. The error rates of DeepFace in gender and race classification is $20.56$ and $38.32$, respectively, whereas FairFace achieves $1.87$ and $19.63$. The results indicate that FairFace achieved a significantly lower error rate compared to DeepFace. Consequently, FairFace is selected as the detector for all subsequent experimental analyses.
@@ 6.1 Cognitive Errors in LLMs
6.2 CoT Effect.
To further understand LLM behaviors, we conduct an error analysis based on their Chain-of-Thought (CoT) Wei et al. (2022) explanations. Our analysis on LLM-generated explanation reveals recurring patterns of cognitive errors: First, representativeness bias is frequently observed. LLMs often rely on overgeneralized group-level assumptions in questions involving race and education, leading to stereotypical justifications such as “White may face fewer systemic barriers” or “Black may face challenges adapting to academic environments.” Second, attribution error emerges in subjective scenarios, such as homelessness, where LLMs infer from isolated factors (e.g., “An Asian male has been homeless for over a decade, suggesting the area aligns with this profile”), thereby projecting specific instances onto broader groups. These findings suggest that LLMs not only produce biased outputs but also mirror human-like cognitive biases in their reasoning, particularly in race- and gender-sensitive contexts.
@@ 7 Related Work
Fairness Issues in Generative AI.
@@
Bias Detection.
With the increasing use of LLMs, bias detection has gained attention. OccuGender Chen et al. (2025) benchmark assesses gender bias in occupational contexts, while Ding et al. (2025) examines cultural and linguistic variations in gender bias. BiasAlert Fan et al. (2024) is a human-knowledge-driven bias detection tool, and Wilson and Caliskan (2024) highlights LLM-induced bias in resume screening, disproportionately affecting black males. BiasAsker Wan et al. (2023) constructs a dataset of 841 groups and 5,021 biased properties. Studies also investigate LLM biases in coding Du et al. (2025), video games Shi et al. (2025), and geolocating tasks Huang et al. (2025b). Bias detection in multimodal models is also emerging. Qiu et al. (2023) investigates gender biases in image captioning metrics, proposing a hybrid evaluation approach. BiasPainter Wang et al. (2024a) is a framework for quantifying social biases by analyzing demographic shifts in generated images. Wan et al. (2024a) provides a comprehensive review of biases in T2I models, identifying mitigation gaps and advocating for human-centered fairness approaches. These studies contribute to improving fairness in generative AI.
Fairness-Accuracy Trade-Off.
@@
Improving Fairness.
To mitigate biases, researchers have proposed various techniques. Jiang et al. (2024) and Shen et al. (2024) improve fairness through fine-tuning and enhanced semantic consistency, while Friedrich et al. (2023) and Li et al. (2025) introduce bias adjustment and fair mapping methods. Su et al. (2024) develops a “flow-guided sampling” approach to reduce bias without modifying model architecture. These methods provide practical strategies for fairness enhancement.
@@ 8 Conclusion
Despite its practical value, Fact-or-Fair still has several limitations that open avenues for future work. (1) Geographic and demographic scope. All 19 statistics are drawn from U.S. datasets and use a coarse set of demographic categories (binary gender and four racial groups). We do not test intersectional identities (e.g., Black women) or protected attributes such as age, disability, religion, or socioeconomic status, so generalising our findings beyond the United States or to richer demographic axes requires caution. (2) Template-based query design. Both objective and subjective queries rely on fixed templates. Although the three cognitive-error contexts increase diversity, they still under-represent the open-ended, multimodal prompts that real users issue. Future work could crowd-source prompts or mine real query logs to improve ecological validity. (3) Reliance on automatic attribute detectors. Image-based evaluations assume perfect gender and race recognition. Even the stronger detector we choose (FairFace) still shows non-negligible error, especially for race ($\approx 20\%$), which can attenuate or inflate fairness scores. These caveats underscore that Fact-or-Fair should be viewed as a configurable testing scaffold rather than a definitive audit. Researchers can extend it with additional regions, demographics, prompts, modalities, and fairness notions to suit their application needs.
@@
We reflect on ethical aspects across data, methodology, potential impact and mitigation: (1) Data provenance and privacy. All 19 indicators (employment, crime, health, etc.) are drawn from U.S. government or inter-governmental releases that are already public, aggregate and anonymized. No personal or proprietary data are used. Because the benchmark relies on coarse categories (Male/Female; Asian/Black/Hispanic/White), it does not enable re-identification of individuals nor infringe on privacy rights. (2) Bias and representational harm. Subjective prompts deliberately surface cognitive errors (representativeness, attribution, in-/out-group bias) to stress-test models. While this can reveal harmful stereotypes, it might also reinforce them when prompts or generated images are taken out of context. We therefore release Fact-or-Fair under a research-only license and accompany it with clear guidance discouraging discriminatory or decision-making use. (3) Downstream misuse. Benchmark scores could be misused to market systems as “fully fair.” To minimize this risk we (i) report both factuality and fairness, (ii) visualize their formal trade-off, and (iii) recommend publishing full score tables rather than single aggregates.
LLM Usage
LLMs were employed in a limited capacity for writing optimization. Specifically, the authors provided their own draft text to the LLM, which in turn suggested improvements such as corrections of grammatical errors, clearer phrasing, and removal of non-academic expressions. LLMs were also used to inspire possible titles for the paper. While the system provided suggestions, the final title was decided and refined by the authors and is not directly taken from any single LLM output. In addition, LLMs were used as coding assistants during the implementation phase. They provided code completion and debugging suggestions, but all final implementations, experimental design, and validation were carried out and verified by the authors. Importantly, LLMs were NOT used for generating research ideas, designing experiments, or searching and reviewing related work. All conceptual contributions and experimental designs were fully conceived and executed by the authors.